OpenStudy (anonymous):

assume a and b are positive integers.decide whether each statements is true or false. (a!)^b=a^b!

OpenStudy (rsadhvika):

try by putting a=3, b=2

OpenStudy (anonymous):

is it true

OpenStudy (rsadhvika):

im not sure you have to try and check

OpenStudy (rsadhvika):

(3!)^2 = 3^2! is this true ?

OpenStudy (anonymous):

i think its tre

OpenStudy (rsadhvika):

3! = 3*2*1 = 6 2! = 2*1 = 2 try putting these values above and check

OpenStudy (rsadhvika):

(3!)^2 = 3^2! 6^2 = 3^2 36 = 9 is this true ?